I need to use the method of substitution to solve these linear equations.
3x-y=3
7x-2y=11

Answer by jim_thompson5910(35256)

Hint:

Solve the first equation for y
3x-y=3

-y=3-3x

y=-3+3x

y=3x-3

Now plug this into the second equation

7x-2y=11

7x-2(3x-3)=11

and solve for x. Then use this to find y.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------